On January 1, 2021, Canseco Plumbing Fixtures purchased equipment for $60,000. Residual value at the end of an estimated four-year service life is expected to be $12,000. The company expects the equipment to operate for 20,000 hours. The equipment operated for 3,700 and 4,500 hours in 2021 and 2022, respectively.

Required:

a. Calculate depreciation expense for 2021 and 2022 using straight-line method.
b. Calculate depreciation expense for 2021 and 2022 using double-declining-balance method.
c. Calculate depreciation expense for 2021 and 2022 using units-of-production using hours operated.

  • Required A

Calculate depreciation expense for 2021 and 2022 using straight-line method.

Straight-Line Depreciation
Choose Numerator: / Choose Denominator: = Annual Depreciation Expense
Formula / = Depreciation Expense
Amounts / = 0
Depreciation Expense
2021
2022
  • Required B

Calculate depreciation expense for 2021 and 2022 using double-declining-balance method.

Double-Declining-Balance Method
Choose Numerator: × Choose Denominator: = Depreciation Expense
Formula Beginning Book Value × = Depreciation Expense
Amount for 2021 × % =
Amount for 2022 × % =
  • Required C

Calculate depreciation expense for 2021 and 2022 using units-of-production using hours operated. (Round "Depreciation per machine hour" answers to 2 decimal places.)

Select formula for Units of Production Depreciation:
Calculate 2021 depreciation expense:
Depreciation per machine hour
Machine hours in 2021
Depreciation in 2021
Calculate 2022 depreciation expense:
Depreciation per machine hour
Machine hours in 2022
Depreciation in 2022
Answer #1

solution-

1. straight line depreciation method

formula for each year depreciation=(cost of machine-salvage value)/estimated useful life of the machine

depreciation expense for

2021=(60000-12000)/4=12000

2022=(60000-12000)/4=12000

2.double declining balance method

formula for depreciation expense=2*straight line depreciation rate*book value at beginning

straight line depreciation rate={deprecation according to straight line method/(cost-salvage value)}*100

={12000/(60000-12000)}*100=25%

year 2021

beginning value=60000

depreciation expense=(2)*(25%)*(60000)=30000

year 2022

beginning value for 2022=beginning value of 2021-depreciation expense for 2021=60000-30000=30000

depreciation expense for 2022=(2)*(25%)*30000=15000

3.units of prodiction method

formula for each year depreciation

=[(cost of machine-salvage value)/expected opertion hours during its useful life]*actual hours operated

depreciation expense for 2021=[(60000-12000)/20000]*3700=8880

depreciation expense for 2022=[(60000-12000)/20000]*4500=10800

note-

depreciaion per machine hour=(cost-salvage value)/total hours expected during the life of machine

for 2021 and 2022=(60000-120000)20000=2.4